against the wind a commercial airline in south america flew 420 miles in 3 hours. with a tailwind the return trip took 2.5 hours. what was the speed of the plane in still air? what was the speed of the wind?
Answer by mananth(16949)

Plane speed =x mph
wind speed =y mph
against wind 3 hours
with wind 2.5 hours

Distance against 420 miles distance with 420 miles
t=d/r against wind -
420.00 / ( x - y )= 3.00
3.00 ( x - y ) = 2.50
3.00 x - 3.00 y = 420.00 ....................1

420.00 / ( x + y )= 2.50
2.50 ( x + y ) = 420.00
2.50 x + 2.50 y = 420.00 ...............2
Multiply (1) by 2.50
Multiply (2) by 3.00
we get
7.50 x + -7.50 y = 1050.00
7.50 x + 7.50 y = 1260.00
15.00 x = 2310.00
/ 15.00
x = 154.00 mph

plug value of x in (1) y
3.00 x -3.00 y = 420.00
462.00 -3.00 -462.00 = 420.00
-3.00 y = 420.00
-3.00 y = -42.00 mph
y = 14.00
Plane speed 154.00 mph
wind speed 14.00 mph

m.ananth@hotmail.ca